# Path Loss Prediction Using Deep Learning

### v1

@@ Abstract
Radio deployments and spectrum planning can benefit from path loss predictions. Obstructions along a communications link are often considered implicitly or through derived metrics such as representative clutter height or total obstruction depth. In this paper, we propose a path-specific path loss prediction method that uses convolutional neural networks to automatically perform feature extraction from high-resolution obstruction height maps. Our methods result in low prediction error in a variety of environments without requiring derived obstruction metrics.
@@ I Introduction
Propagation models are used by the wireless industry and regulatory bodies to aid in analysis to enhance wireless coverage and avoid unwanted interference. Path loss modeling, a common approach to propagation modeling, aims to predict point-to-point (P2P) wireless signal attenuation along the communications link from the transmitter (Tx) to the receiver (Rx). The three main factors in predicting path loss are carrier frequency, link distance, and obstructions in the line of sight (LOS). In LOS conditions, free space path loss (FSPL), which uses frequency and distance, is sufficient for accurate P2P path loss predictions. In non-LOS conditions, however, FSPL tends to severely underestimate path loss.
@@
LOS obstructions are obtained through surface-level variations, which are captured by the geographic information system (GIS) quantity digital surface model (DSM). DSM is measured relative to sea level and captures all heights, which can include ground-level variations (terrain) as well as trees, buildings, and other structures (clutter). Where no clutter is present, DSM models the terrain.
Traditional path loss models, such as Longley-Rice [1], explicitly account for terrain information but do not explicitly consider the presence of clutter. Consequently, these predictions can underestimate path loss in links dominated by clutter.
@@
Modern path loss models that consider clutter information have been shown to significantly decrease prediction error in non-LOS links. One such model, P.1812-6 [2] (recommended by the International Telecommunications Union Radiocommunication Sector (ITU-R)), considers clutter along the link in the form of representative clutter heights. The values of these heights are computed using local information regarding the types of clutter present. This yields more accurate predictions, but there is room for improvement due to the approximations imposed by the representative clutter height computation and the emphasis on diffraction losses [3].
@@
Previous work in [3] shows improvement over P.1812-6 in non-LOS conditions, by capturing GIS information in one novel feature: total obstruction depth. This scalar feature is computed using surface information along the direct path. Alongside frequency and distance, obstruction depth yields a root mean square error (RMSE) in the range of 6-8 dB [3] as compared to 8-13 dB for P.1812-6. While using total obstruction depth as a feature can produce lower errors, the feature is a single scalar value derived from the surface profile. It does not explicitly consider the surface height profile, which can influence path loss in more complicated ways due to diffraction and Fresnel zone effects [4].
@@
The goal of this paper is to explicitly utilize the surface profile to train a path-specific propagation model. Leveraging high-resolution clutter maps and deep neural networks to directly predict path loss, we aim to study whether high-dimensional spatial features can be extracted more reliably by neural networks than by human-defined feature engineering. The secondary goal is to compare whether these complex features are improved by adding the transverse dimension (the dimension perpendicular to the link distance).
Previous work [5] uses map-based inputs to train path loss models limited to two frequencies within a single area, with no geographically distinct tests. Random cross-validation using spatial data can lead to data leakage which can lead to overfitting [6]. Our work aims to construct generalized models that can predict path loss in unseen scenarios, which can only be judged through blind testing.
@@
This paper describes path loss models trained on one- and two-dimensional (1D & 2D) surface height data alongside measurement data. Section II describes the pre-processed datasets used to train and evaluate the path loss models. Section III justifies the multi-dimensional features used from the data and details the model architectures used for the 1D and 2D clutter height data. Section IV describes the training and evaluation process and results. Finally, Section V summarizes our findings and presents potential future directions.
@@ II-A Dataset Description
There are two types of data used in this work: radio measurement data, and geographic elevation/clutter data. For radio measurement data, we use drive test measurement data from the United Kingdom Office of Communications (UK Ofcom) [7], which can be accessed by the general public through Ofcom’s open data repository [8]. This dataset was collected between 2015 and 2018, and provides 8.2 million measurements across six frequencies (449, 915, 1802, 2695, 3602, and 5850 MHz) in seven cities in the UK. Each city uses a single Tx with varying Rx locations. Link distances vary from 1 m to 78 km, with a median of 8 km.
For clutter data, we use the UK open digital elevation model (DEM) [9]. This dataset provides DSM at a 1 m resolution for various areas across the UK. DEM data is available for six out of seven drive test cities: Boston, London, Merthyr Tydfil, Nottingham, Southampton, and Stevenage. We use data from these six cities to train and test our path loss models.
The following subsection discusses the steps taken to filter, pre-process, and join these two datasets.
@@ II-B1 Ofcom drive test data
Firstly, we use measurements with received signal level ($RSL$) above the system noise floor provided in the dataset, by a 6 dB margin, as shown in (1a). We exclude measurements near and below the noise floor to ensure the model is trained on reliable data. Similarly, we use only links with a total distance $d$ above 50 meters as in (1b):
@@
This allows training and evaluation to be conducted only on samples that have a reasonable expectation (and benefit) to be modeled. After both filtering steps, 4.2 million measurements remain for joining with DSM data.
@@ II-B2 Digital surface model data
Finally, we apply earth curvature correction [10, eq. (7-16)] to the DSM heights relative to the Tx, using a spherical radius of 6364.75 km. The 1D path profiles are then cropped from the direct (center) path of the 2D path profiles, yielding an array of shape $(\left\lfloor{d}\right\rfloor\times 1)$ for each measurement.
@@ III Feature and Architecture Selection
We formulate this work as an image problem, where each input feature forms a channel [11] of an image (1D or 2D).
@@ III-A Feature Selection
Before computing the feature channels, images are re-sampled to have a constant direct path length of 256 pixels. While the majority of path profiles are down-sampled, some of the shorter links result in up-sampling to 256. Initial experimentation was conducted with training on variably-sized arrays with lengths proportional to the corresponding link distances. This approach yielded poor results, achieving 8-10 dB RMSE with significantly more computational requirements. Re-sampling to a uniform size allows for the design of a single architecture that generalizes well for all link distances.
In the re-sampled arrays, we include the essential features, frequency and distance, as their own channels. Frequency (MHz) is encoded as a channel with a constant value, while the distance channel is generated by computing the 2D Euclidean distance in metres from the Tx to each pixel. The distance channel serves the dual purpose of encoding relative spatial locations [12] while providing the non-re-sampled distances.
We use two additional channels to capture path profile information: (1) surface and (2) direct path. Surface (DSM) provides the model with the path profile height information and direct path provides the antenna locations, heights and the direct path from Tx to Rx antennas. In the case of the 2D path profiles, the pixels in the center of the transverse dimension contain the direct path heights and are 0 elsewhere in the channel.
The path profile extraction process and results are shown in Fig. 1 for a link in London with a distance of 189 m. The final dataset contains samples of shape: $\ (4~{}\mathrm{channels}\times 256~{}\mathrm{length}\times N~{}\mathrm{width})$, where $N$ is 1 for the 1D method and 61 for the 2D method.
@@
Surface Normalized (0, 1) using heights of surface and direct path channels Per sample
@@
Re-sampled and normalized inputs to the 1D and 2D path loss models are shown in Fig. 2 and Fig. 3 respectively for a 449 MHz link in London.
Figure 2: Re-sampled and normalized four-channel input to 1D CNN.
Figure 3: Re-sampled and normalized surface channel input to 2D CNN.
III-B Model Architectures
Given the spatial structure of the data, we employ convolutional neural networks (CNNs) [11], treating the data as images. We design two CNNs: one for processing the 1D path profiles, and another for the 2D path profiles. The 2D CNN is designed identically to the 1D CNN, except all kernels are square windows of the same size, rather than linear. Each convolution layer has a kernel size of 3, stride size of 2, and zero-padding size of 1. Each max pool layer has a kernel size of 3, stride size of 1, and “same” zero-padding [11].
During training, we use dropout at a rate of 25% on the flattened features to minimize overfitting. The extracted features are input to a fully connected network (FCN) [11] with 256 hidden units and rectified linear unit (ReLU) activation. The final output is path loss.
We design the 2D CNN model identically to the 1D model on the length axis. The transverse axis kernel sizes are chosen such that the image dimensions decrease at the same rate, resulting in the same number of features. Given the low RMSE in [3] with only 3 features, 4096 features are likely to be sufficient, regardless of the feature extraction process.
TABLE II: Structure of 1D and 2D CNNs
@@ III-B Model Architectures
(channels $\times$ length $\times$ [1D or 2D width])
Input 4 $\times$ 256 $\times$ [1 or 61]
Convolution + ReLU 32 $\times$ 256 $\times$ [1 or 61]
Max pool 32 $\times$ 128 $\times$ [1 or 31]
Convolution + ReLU 64 $\times$ 128 $\times$ [1 or 31]
Max pool 64 $\times$ 64 $\times$ [1 or 16]
Convolution + ReLU 128 $\times$ 64 $\times$ [1 or 16]
Max pool 128 $\times$ 32 $\times$ [1 or 8]
Convolution + ReLU 256 $\times$ 32 $\times$ [1 or 8]
Max pool 256 $\times$ 16 $\times$ [1 or 4]
Convolution + ReLU 512 $\times$ 16 $\times$ [1 or 4]
Max pool 512 $\times$ 8 $\times$ [1 or 2]
Convolution + ReLU 1024 $\times$ 8 $\times$ [1 or 2]
Max pool 1024 $\times$ 4 $\times$ 1
@@
The network structure is summarized in Table II. The 1D and 2D models have total parameter counts of 3 146 593 and 7 337 569 respectively. Both neural networks are developed, trained, and evaluated in PyTorch [13]. The following section describes the training regimen and evaluation results.
@@ IV Training and Results
To decrease training times, a random subset of the filtered data is taken to compare the two models. 10 000 samples are taken from each frequency/city combination, resulting in 60 000 samples per city, for a total of 360 000 samples.
We perform a six-city cross-validation, where each city is separately held out as the test set, with the remaining five cities split between training and validation. We use adaptive moment estimation (Adam) [14] with a learning rate of 0.0001 and batch size of 256 for 200 epochs. The model with the lowest validation RMSE is used to evaluate test RMSE.
@@ IV-A Validation Split
Deep learning models are more likely to overfit to GIS arrays (feature-rich) than scalar features (feature-sparse) for the same amount of data, given the difference in dimensionality [15]. To mitigate this, for each city holdout, we geographically split each of the remaining five cities using 80% for training and 20% for validation. In each of these five cities, an initial angle $\theta$ is chosen. Rotating counter-clockwise about the Tx, the first 20% of Rx locations encountered are used as validation samples, with the remaining 80% used for training. A geographic split for each of the five non-holdout cities provides diverse yet geographically distinct training and validation sets. This split is necessary with this feature set, as initial studies using random splits caused significant overfitting due to high correlation with training sets. Fig. 4 shows an example of a training-validation split in London.
Figure 4: Example of geographic training-validation split in London.
@@ IV-B Evaluation of Result Consistency
Table III summarizes the cross-validation results with a mean and standard deviation (SD) of RMSE for both models. We also compare the mean results with ITU-R P.1812-6 [2] (the industry standard for 1-6 GHz, hereafter referred to as P.1812) and a three-feature FCN using frequency, distance, and total obstruction depth [3].
TABLE III: Cross-Validation Holdout RMSE (dB) Across 10 Runs
Holdout 1D CNN 2D CNN Benchmarks [3]
Mean SD Mean SD P.1812 FCN
Boston 7.78 0.60 7.18 0.27 11.4 7.56
London 9.95 0.89 7.75 0.47 8.8 7.44
Merthyr Tydfil 8.09 0.44 8.07 0.39 13.4 7.22
Nottingham 7.16 0.30 6.99 0.23 12.6 7.07
Southampton 6.62 0.45 6.35 0.42 9.5 6.64
Stevenage 7.65 0.31 7.73 0.21 12.3 8.81
Mean 7.88 0.50 7.35 0.33 11.3 7.46
Figure 5: Histogram of prediction errors for first London holdout.
The 2D models generally have lower RMSE than 1D models, benefiting from the additional transverse information in the 2D path profiles, and resulting in more pertinent features.
@@
Compared to the FCN, our 2D model shows lower mean RMSE on four of six holdouts, with higher mean RMSE on the other two holdouts (London and Merthyr Tydfil). Our 1D model has higher mean RMSE than the FCN on four of six holdouts. Although some of the holdouts show lower RMSE than the FCN, the fact that other holdouts have higher RMSE in the presence of richer feature sets suggests a further need for model refinement and feature selection.
Generally, our modeling approaches provide significant improvement over the current industry standard. Across all models and holdouts, our CNNs have lower RMSE than P.1812 in 11 of 12 cases, with the only exception being the 1D London-holdout model. This indicates that higher-dimensional information outperforms representative clutter heights for predictive modeling purposes. Fig. 5 shows the error distribution for the CNNs and P.1812 for a single-run London holdout (the lowest drive test RMSE for P.1812). Across runs and holdouts, the 2D model consistently has mean error closest to 0 dB.
Merthyr Tydfil and London are the two holdouts with the highest RMSE. Despite sequential overfitting [16] to the London holdout, the other four cities show strong generalization. Since London contains the highest clutter density among the six cities, and Merthyr Tydfil contains the highest terrain density, it is more challenging to generalize high-dimensional GIS features to these two holdouts. Test performance may be improved by providing the difference between terrain and clutter via a terrain channel.
@@
Future studies will investigate additional feature channels to decrease test RMSE, such as height scaling and Fresnel information within the path profiles, terrain to distinguish between clutter and ground, and type of clutter. Future work may also investigate further validation splitting techniques to minimize overfitting.

### v2

@@ Abstract
Radio deployments and spectrum planning benefit from path loss predictions. Obstructions along a communications link are often considered implicitly or through derived metrics such as representative clutter height or total obstruction depth. In this paper, we propose a path-specific path loss prediction method that uses convolutional neural networks to automatically perform feature extraction from high-resolution obstruction height maps. Our methods result in low prediction error in a variety of environments without requiring derived metrics.
@@ I Introduction
Path loss modeling is a tool used by the wireless industry and regulatory bodies to optimize wireless coverage and avoid unwanted interference. Path loss models predict point-to-point (P2P) wireless signal attenuation along the communications link from the transmitter (Tx) to the receiver (Rx). Path loss is primarily affected by frequency, link distance, and obstructions in the line of sight (LOS). In LOS conditions, free space path loss (FSPL), which uses frequency and distance, is sufficient for accurate P2P path loss predictions. In non-LOS conditions, however, FSPL typically underestimates path loss.
@@
LOS obstructions are identified from surface-level variations, captured by the geographic information system (GIS) quantity digital surface model (DSM). DSM captures heights relative to sea level: ground-level variations (terrain) as well as trees, buildings, and other structures (clutter). Where there is no clutter, DSM models the terrain.
@@
Traditional path loss models, such as Longley-Rice [1], explicitly account for terrain information but do not explicitly consider the presence of clutter. Consequently, these predictions can underestimate path loss in links dominated by clutter. Conversely, modern path loss models that consider clutter information have significantly lower prediction errors in non-LOS links. One such model, P.1812-6 [2] (recommended by the International Telecommunications Union Radiocommunication Sector (ITU-R)), considers clutter along the link in the form of representative clutter heights. The heights are calculated based on local clutter types, improving prediction accuracy but remaining limited by approximations in clutter height estimation and focusing on diffraction losses [3].
@@
Previous work in [3] shows improvement over P.1812-6 in non-LOS conditions, by capturing GIS information in one novel feature: total obstruction depth. This scalar feature is computed using surface information along the direct path. Alongside frequency and distance, obstruction depth yields a root mean square error (RMSE) in the range of 6-8 dB as compared to 8-13 dB for P.1812-6 [3]. While using total obstruction depth as a feature can produce lower errors, it does not explicitly consider the surface height profile, which can influence path loss in more complicated ways due to diffraction and Fresnel zone effects [4].
@@
The goal of this paper is to explicitly utilize the surface height profile to train a path-specific propagation model. Leveraging high-resolution clutter maps and deep neural networks to directly predict path loss, we aim to study whether high-dimensional spatial features can be extracted more reliably by neural networks than by human-defined feature engineering.
Previous work includes [5, 6, 7], which demonstrate generalized path loss prediction in spatial holdouts, using simulated data in [5] or measurement data in [6, 7]. Conducting geographic cross-validation within a single region, [8] provides a robust assessment of generalization using measurement data. Similarly, [9] uses map-based inputs to train path loss models limited to two frequencies in a single region. Our methods use geographic cross-validation across multiple cities–a prerequisite to generalized models that can accurately predict in new regions [10]. Our work differs from [5, 6] by training and testing exclusively on measurements, from [3, 5, 6] by explicitly using the surface height profile as an input, and from [5, 6, 7, 8, 9], by conducting geographic cross-validation across multiple cities. To our knowledge, the novel combination of explicitly using the surface height profile, exclusively training and testing on measurements, and evaluating with geographic cross-validation across multiple cities has not been explored.
@@
This paper describes path loss models trained on 2-D surface height data alongside measurement data. Section II describes the pre-processed datasets used to train and evaluate the path loss models. Section III explains the multi-dimensional features extracted from the data and the model architecture. Section IV describes the training and evaluation process and results. Finally, Section V summarizes our findings.
@@ II-A Dataset Description
In this paper, we use two datasets: radio measurement data, and GIS elevation/clutter data. For radio measurement data, we use drive test measurement data from the United Kingdom Office of Communications (UK Ofcom) [11], which can be accessed by the general public through Ofcom’s open data repository [12]. This dataset was collected between 2015 and 2018, and provides 8.2 million measurements across six frequency bands (449, 915, 1802, 2695, 3602, and 5850 MHz) in seven cities in the UK. Each city uses a single Tx with varying Rx locations. Link distances vary from 1 m to 78 km, with a median of 8 km.
For clutter data, we use the UK open digital elevation model (DEM) [13]. This dataset provides DSM at a 1 m resolution for various areas across the UK in 2021. DEM data is available for six out of seven drive test cities: Boston, London, Merthyr Tydfil, Nottingham, Southampton, and Stevenage. We use data from these six cities to train and test our path loss models.
@@ II-B1 Ofcom drive test data
Firstly, we use measurements with received signal level ($RSL$) above the system noise floor provided in the dataset, by a 6 dB margin, as shown in (1a). Similarly, we use only measurements with a link distance $d$ above 50 meters as in (1b):
@@
This allows training and evaluation to be conducted only on samples that have a reasonable expectation (and benefit) to be modeled. After both filtering steps, 4.2 million measurements remain for joining with DSM data. Link distances in this subset vary from 50 m to 57 km, with a median of 7 km.
@@ III Feature and Architecture Selection
We formulate this work as an image problem, where each input feature forms a channel [16] of a 2-D image.
@@ III-A Feature Selection
Before computing the feature channels, we re-sample all images to have a direct path length of 256 pixels, using bilinear interpolation [17] . While the majority of path profiles are down-sampled, some of the shorter links result in up-sampling to 256. Initial experimentation with variably-sized array lengths proportional to the corresponding link distances yielded poor results with significantly more computational requirements. Re-sampling to a uniform size allows for the design of a single architecture that generalizes well for all link distances. We choose a resolution of 256 pixels for its consistently low validation RMSE during architecture selection, with higher resolutions suffering from overfitting.
In the re-sampled arrays, we include the essential features, frequency, and distance, as their own channels. Frequency [MHz] is encoded as a channel with a constant value, while the distance channel is generated by computing the 2-D Euclidean distance [m] from the Tx to each pixel. The distance channel serves the dual purpose of encoding relative spatial locations [18] while providing the non-re-sampled distances.
We use two additional channels to capture path profile information: (1) DSM and (2) direct path. DSM provides the path profile heights and direct path provides the antenna locations and heights, and the 3D path from Tx to Rx. Antenna heights relative to sea level are obtained by adding the heights provided by Ofcom to the terrain. This channel contains the direct path heights in the center of the transverse dimension and 0 elsewhere.
Fig. 1 shows the path profile extraction process and result for a link in London with a distance of 411 m. The final dataset contains samples of shape: $\ (4~{}\mathrm{channels}\times 256~{}\mathrm{length}\times W~{}\mathrm{width})$, where $W=61$.
@@
Surface Normalized to [0, 1] using heights of surface and direct path Per sample
@@
Fig. 2 shows a re-sampled and normalized input to the path loss model for the same link shown in Fig. 1. We also compare the effect of re-sampling on a typical link, showing both re-sampled and full DSM values for the center (direct path) of an 8 km in Fig. 3. Features such as thin building peaks are filtered out, lowering precision. However, the aforementioned validation improvements suggest that re-sampling provides a regularization effect.
Figure 2: Re-sampled and normalized surface channel input to CNN.
Figure 3: Full size and re-sampled center of path profile for a link in London.
III-B Model Architecture
Given the spatial structure of the data, we employ convolutional neural networks (CNNs) [16], treating the data as images. Each convolution layer has a kernel size of 3$\times$3, stride size of 2$\times$2, and zero-padding size of 1$\times$1. Each max pool layer has a kernel size of 3$\times$3, stride size of 1$\times$1, and “same” zero-padding [16].
During training, we use dropout at a rate of 25% on the 4096 flattened features to minimize overfitting. The extracted features are input to a fully connected network (FCN) [16] with 256 hidden units and rectified linear unit (ReLU) activation. The final output is path loss. Given the low RMSE in [3] with only 3 features, 4096 features are likely to be sufficient, regardless of the feature extraction process.
TABLE II: Summary of CNN Architecture
@@ III-B Model Architecture
(channels $\times$ length $\times$ width)
Input 4 $\times$ 256 $\times$ 61
Convolution + ReLU $\xrightarrow{}$ Max pool 32 $\times$ 256 $\times$ 61 $\xrightarrow{}$ 32 $\times$ 128 $\times$ 31
Convolution + ReLU $\xrightarrow{}$ Max pool 64 $\times$ 128 $\times$ 31 $\xrightarrow{}$ 64 $\times$ 64 $\times$ 16
Convolution + ReLU $\xrightarrow{}$ Max pool 128 $\times$ 64 $\times$ 16 $\xrightarrow{}$ 128 $\times$ 32 $\times$ 8
Convolution + ReLU $\xrightarrow{}$ Max pool 256 $\times$ 32 $\times$ 8 $\xrightarrow{}$ 256 $\times$ 16 $\times$ 4
Convolution + ReLU $\xrightarrow{}$ Max pool 512 $\times$ 16 $\times$ 4 $\xrightarrow{}$ 512 $\times$ 8 $\times$ 2
Convolution + ReLU $\xrightarrow{}$ Max pool 1024 $\times$ 8 $\times$ 2 $\xrightarrow{}$ 1024 $\times$ 4 $\times$ 1
@@
The network structure is summarized in Table II. The model has 7 337 569 total parameters. Our CNNs are developed, trained, and evaluated in PyTorch [19]. The following section describes the training regimen and evaluation results.
@@ IV Training and Results
We randomly sample 10 000 unique links from each frequency/city combination, resulting in 60 000 samples per city, for a total of 360 000 samples. This ensures a uniform training distribution across different cities and frequency bands.
We use a six-city cross-validation, where each city is held out as the test set, with the remaining five cities split between train and validate. We use adaptive moment estimation (Adam) [20], mean squared error loss, a learning rate of 0.0001, and a batch size of 256 for 200 epochs. The model with the lowest validation RMSE is used to evaluate test RMSE for each run.
@@ IV-A Validation Split
Deep learning models are more likely to overfit to GIS arrays (feature-rich) than scalar features (feature-sparse) for the same amount of data, given the difference in dimensionality [21]. To mitigate this, for each city holdout, we geographically split each of the remaining five cities using 80% for training and 20% for validation. In each of these five cities, an initial angle $\theta$ is chosen. Rotating counter-clockwise about the Tx, the first 20% of Rx locations encountered are used as validation samples, with the remaining 80% used for training. A geographic split for each of the five non-holdout cities provides diverse yet geographically distinct training and validation sets. This split is necessary with this feature set, as initial studies using random splits caused significant overfitting due to high correlation with training sets.
@@ IV-B Evaluation of Result Consistency
Table III summarizes the cross-validation results with the mean and standard deviation (SD) of test RMSE for each holdout. We also compare the mean results with ITU-R P.1812-6 [2] (the industry standard for 1-6 GHz, hereafter referred to as P.1812) and a three-feature FCN using frequency, distance, and total obstruction depth [3].
TABLE III: Cross-Validation Test RMSE [dB] Across 10 Runs
CNN Benchmarks [3]
Holdout Mean SD P.1812 FCN
Boston 7.18 0.27 11.4 7.56
London 7.75 0.47 8.8 7.44
Merthyr Tydfil 8.07 0.39 13.4 7.22
Nottingham 6.99 0.23 12.6 7.07
Southampton 6.35 0.42 9.5 6.64
Stevenage 7.73 0.21 12.3 8.81
Mean 7.35 0.33 11.3 7.46
The model generally produces lower RMSE for lower frequencies and shorter distances, as shown in Table IV. The major errors occur on the extremely long links. Across all runs and holdouts, the CNN achieves 7.37 dB RMSE on the shortest 99.9% of unique holdout links ($<$45 km), and 12.04 dB RMSE on the remaining 0.1% ($<$800 links). This implies that re-sampling has a significant effect on performance, especially for longer links. Future work should ensure more representation of long links in the training set, or only operate the model on shorter links to ensure confidence in prediction.
TABLE IV: Subset RMSEs Computed Across All Runs
Subset Holdout RMSE [dB]
449 MHz 6.31
915 MHz 7.10
1802 MHz 7.24
2695 MHz 7.32
3602 MHz 7.99
5850 MHz 8.15
*Shortest 99.9% 7.37
Longest 0.1% 12.04
*99.9% of links are below 45 km.
@@
Compared to the FCN, our model shows lower mean RMSE on four of six holdouts, with higher mean RMSE on the other two holdouts (London and Merthyr Tydfil). Although four of six holdouts show lower RMSE than the FCN, the fact that other holdouts have higher RMSE in the presence of richer feature sets suggests a further need for model refinement and feature selection. Though these trends may differ when attempting to generalize to other countries and measurements, using multi-city cross-validation suggests that the result would likely generalize to other unseen regions.
Our methods simplify modeling by eliminating feature engineering but require a larger initial investment in training time–80 minutes per run vs. 2 minutes for the FCN. The inference speed of the CNN and the FCN are 4000 and 10 000 samples/s, respectively, with the same hardware (Amazon Web Services g5.4xlarge), which are both computationally feasible.
Our CNN-based approach provides significant improvement over the current industry standard, achieving lower RMSE than P.1812 in all six holdouts. This implies that higher-dimensional information outperforms representative clutter heights in predictive models.
Merthyr Tydfil, London, and Stevenage are the three highest holdout RMSEs. The London holdout had consistently high RMSE while testing several hyperparameter configurations. To mitigate this, we tuned our hyperparameters to minimize London RMSE (sequential overfitting [22]). Choosing a hyperparameter configuration that brought London under 8 dB RMSE resulted in a trade-off where Stevenage RMSE was increased to a similar value to London. Despite this, London is still in the upper half of holdout RMSEs, with the other half showing strong generalization. Since London contains the highest clutter density among the six cities, and Merthyr Tydfil contains the highest terrain density, it is more challenging to generalize high-dimensional GIS features to these two holdouts. Stevenage is also difficult to generalize to because it is a relatively new suburb, with newer building materials and different electromagnetic behavior. Additionally, there is likely more rapid expansion occurring compared to the other holdouts, potentially making the three-year discrepancy between measurement and DSM data more significant. Test performance in all three holdouts may be improved with additional channels, such as a terrain channel to differentiate between terrain and clutter, or a building materials channel. Aligning the datasets in time is also likely to decrease RMSE.
@@
Future studies will investigate additional feature channels to decrease test RMSE, such as height scaling and Fresnel information within the path profiles, terrain to distinguish between clutter and ground, and type of clutter. Future work may also study further validation splitting techniques to minimize overfitting. Finally, including DSM height information in the area behind the Tx and Rx antennas will likely reduce RMSE as a result of its impact on path loss.